### v1

# Literature Review, Methodology, Software, Data curation, Implementation, Experimental analysis, Writing - Original draft preparation

Abstract
Phishing email is a serious cyber threat that tries to deceive users by sending false emails with the intention of stealing confidential information or causing financial harm. Attackers, often posing as trustworthy entities, exploit technological advancements and sophistication to make detection and prevention of phishing more challenging. Despite extensive academic research, phishing detection remains an ongoing and formidable challenge in the cybersecurity landscape. Large Language Models (LLMs) and Masked Language Models (MLMs) possess immense potential to offer innovative solutions to address long-standing challenges. In this research paper, we present an optimized, fine-tuned transformer-based DistilBERT model designed for the detection of phishing emails. In the detection process, we work with a phishing email dataset and utilize the preprocessing techniques to clean and solve the imbalance class issues. Through our experiments, we found that our model effectively achieves high accuracy, demonstrating its capability to perform well. Finally, we demonstrate our fine-tuned model using Explainable-AI (XAI) techniques such as Local Interpretable Model-Agnostic Explanations (LIME) and Transformer Interpret to explain how our model makes predictions in the context of text classification for phishing emails.
keywords:
[…]
1 Introduction
The rapid advancement of internet technology has significantly changed the way people communicate online while also raising more security issues (Salloum et al., 2021). One of the prevalent security concerns, the cyberattack termed “email phishing” poses grave risks such as identity or login credentials theft, financial repercussions, and harm to an organization’s reputation. In this current decade, email phishing has changed throughout time to become a more effective method for carrying out fraudulent actions. As phishers continuously innovate new techniques, detecting phishing has become increasingly challenging, posing a formidable task for law enforcement agencies, governments, and organizations combating this cyber threat. To evade falling victim to this threat, organizations need to adeptly detect and prevent these attacks at scale by adopting effective phishing detection techniques. Leveraging Artificial Intelligence (AI) techniques like Machine Learning, Deep Learning, and Hybrid Learning which are already utilized for phishing attack detection, can significantly bolster these efforts (Basit et al., 2021) (Sarker, 2023). Moreover, because of the rapid advancements in technology, new methods based on AI are being developed daily in the cybersecurity field.
Transformer-based models become more popular in the context of text classification. The transformer is an innovative neural design that uses the attention mechanism to encode incoming data into powerful characteristics (Han et al., 2021). Transformer-based models have greatly improved the processing, comprehension, and interpretation of text data inputs, which has had a revolutionary effect on the development of phishing classification techniques (Jamal et al., 2023). Large Language Models like OpenAI’s ChatGPT (Generative Pre-trained Transformer) and Bard have transformed the landscape of natural language creation and comprehension, capturing considerable attention from the general public (Zhao et al., 2023; Yao et al., 2023). The increasing complexity of LLMs is radically changing artificial intelligence and has the potential to transform algorithm development and application by understanding human language through extensive text data patterns and correlations.
Many transformer-based models used in natural language processing (NLP), including well-known ones like OpenAI’s GPT and Google’s BERT (Bidirectional Encoder Representations from Transformers), are categorized as Large Language Models because of their ability to comprehend and produce text that is similar to that of a human. BERT model, a type of transformer-based neural network architecture pre-trained on extensive text corpora to grasp contextual word embeddings (Koroteev, 2021), become increasingly efficient and valuable in text classification and generation, including tasks like spam and phishing detection, sentiment analysis (Singh et al., 2021), and more other fields. BERT is designed to take into account both left and right contexts at every level to pre-train complex bidirectional representations from unlabeled text (Devlin et al., 2018). With simply one extra output layer, this method allows the pre-trained BERT model to be refined, leading to the development of state-of-the-art models for a variety of applications. In order to model language, BERT uses a Masked Language Model, in which certain words in a sentence are randomly replaced with a unique token (usually [MASK]). Next, by examining the contextual clues offered by nearby words, the model is trained to infer the original words. BERT makes use of a Transformer architecture technique called “self-attention”, which allows it to recognize inter-word relationships in text sequences and assess the relative importance of individual words within a sentence in order to understand the context. Based on the transformer layers number and parameters BERT model has different variants such as: BERT-base (Khadhraoui et al., 2022), ALBERT (Lan et al., 2019), RoBERTa (Liu et al., 2019), DistilBERT (Sanh et al., 2019), etc.
Despite BERT’s versatility across NLP tasks, comprehending its internal mechanisms often poses challenges, resembling a “black box” due to its complexity. Unraveling how the model operates and forms predictions demands considerable effort (Xu et al., 2019). XAI techniques emerge as a vital solution, offering significant perceptions of the model’s reasoning and illuminating the decision-making process (Hoffman et al., 2018; Anan et al., 2023). Recently, Sarker (Sarker, 2024) has also highlighted the significance of explainability analysis in various perspectives in the context of cybersecurity, in which we are interested to explore more with experiments. XAI techniques can illuminate the words or features deemed most influential by the BERT model in specific text classification tasks of any nature. In this research, our main goal is to show transparency in phishing email detection with explainable AI. The major contribution of the work is as follows:
  • We present a transformer-based Phishing email detection model, a fine-tuned version of DistilBERT.
  • Two different XAI techniques LIME and Transformer Interpret are utilized to explore the interpretability of the proposed fine-tuned model and visualize the scoring percentage of those techniques on prediction.
  • We demonstrate our model performance on both imbalanced and balanced datasets.
The remaining sections of the paper are structured as follows: Section 2 delves into related research works. Section 3 offers a comprehensive explanation of the proposed model’s framework and methodology. In Section 4, the experimental outcomes and results are elaborated upon. Section 5 examines the transparency of the model using XAI. Moreover, within Section 6, there is an extensive discussion of the results. Lastly, Section 7 presents the concluding remarks and outlines the future prospects of this work.
2 Literature Review
The field of cybersecurity has grown increasingly important and complex in the modern world. There are several subfields within cybersecurity, including data security, application security, network security, etc. Cyber-attacks of all kinds are becoming more and more frequent, and they can destroy or steal imported data. In addition, in this cybersecurity field, there are an enormous amount of new and creative techniques being developed daily for identifying and preventing cyberattacks. For the improvement of phishing email detection and prevention systems, many researchers have developed a wide range of novel strategies to counteract phishing emails, utilizing state-of-the-art methods from deep learning, machine learning, and hybrid models. Phishing is still a serious cybersecurity concern and one of the main ways hackers breach networks.
Detecting phishing attempts early on is critical for businesses nowadays. The security of an organization can be significantly increased by machine learning, which can identify potential assaults early on (Apruzzese et al., 2023). In ref. (Yasin and Abuhasan, 2016), a text processing, data mining, and knowledge discovery approach was combined to create an intelligent phishing email detection model. They utilized text stemming and WordNet ontology in pre-processing to enrich the model with synonyms and employed five machine learning classification algorithms (J48, Naïve Bayes, Support Vector Machine (SVM), Multi-Layer Perceptron (MLP), and Random Forest (RF)). They have notably achieved 99.1% accuracy with RF and 98.4% using J48 for an accredited dataset. The study (Harikrishnan et al., 2018) employed techniques such as SVD and NMF for feature extraction and dimensionality reduction from legitimate and phishing email text data. For the numeric representation of phishing emails, the author integrated TFIDF with the SVD and NMF, and Decision Trees, Logistic Regression, Random Forest, Naive Bayes, K-Nearest Neighbors (KNN), AdaBoost, and SVM have been utilized as classifiers. Hamid et al. (2013) presented a hybrid feature selection technique combining content-based and behaviour-based approaches to detect phishing emails. Identifying the behaviour-based features in phishing emails was the main objective of this research because an attacker cannot disguise these types of features. They examined the overall process with a phishing dataset that consists of 6923 emails where phishing emails were collected from the Nazario dataset and legitimate emails from the SpamAssassin dataset and achieved 94% accuracy with this hybrid feature selection technique. With an emphasis on the importance of top-ranking features, the study (Zamir et al., 2020) examines feature selection methods (GR, IG, Relief-F, RFE) on a dataset of 11,055 phishing sites with 32 attributes. To improve classifier accuracy, it suggests new attributes that are generated from the strongest and weakest features. A variety of supervised learning methods with differing feature sets are compared, such as RF, SVM, bagging, kNN, Neural Network (NN), and j48. The proposed method using RF showed 97.3% and using stacking achieved 97.4% accuracy.
In cybersecurity, deep learning has great potential, particularly for detecting network intrusions, classifying malware, and recognizing phishing attempts.Alhogail and Alsabih (2021) proposed a deep learning-based phishing email detection model based on graph convolutional network (GCN) and natural language processing (NLP). A publicly available balanced phishing dataset has been utilized where the authors used the email body’s text to extract features that improved the proposed model’s accuracy. The model showed 98.2% accuracy with a low false-positive rate of 0.015%. For phishing detection, a new Intelligent Cuckoo Search (CS) Optimization Algorithm with a Deep Learning-based Phishing Email Detection and Classification (ICSOA-DLPEC) approach has been proposed in (Brindha et al., 2023). Data preprocessing (email cleaning, tokenization, and stop-word elimination) techniques were employed to derive the important features from the Enron email dataset. The Gated Recurrent Unit (GRU) model was used with a CS algorithm for the fine-tuning of the parameter of the proposed model and achieved a high accuracy of 99.72%.Dewis and Viana (2022) employed Long short-term memory (LSTM) model with NLP to detect both phishing and spam email detection. They developed a phish responder which is a Python-based command line solution and used deep learning and NLP for this classification task. The proposed LSTM model achieved 99% accuracy for the text-based dataset. Fang et al. (2019) introduced THEMIS, a phishing email detection model built on an improved recurrent convolutional neural networks (RCNN) model which was improved by using the Bidirectional Long Short-Term Memory (Bi-LSTM). 99.84% accuracy is demonstrated by the suggested THEMIS model on datasets that are not balanced where the positive rate (FPR) is 0.043%. Other deep learning algorithms, such the reinforcement learning, Recurrent Neural Network (RNN), and Deep Belief Network (DBN), are also in use (Zhang and Li, 2017; Bahnsen et al., 2017; Smadi et al., 2018).
Moreover, transformer-based models are also making a good impact in the cybersecurity field.Thapa et al. (2023) concentrated on two specific models: RNN and BERT to test how well these models could identify phishing emails. They looked at how well these models performed when the email data was distributed unevenly across several organizations and when added more data. In this experiment, the BERT model’s test accuracy was 96.1%. In addition, another study (Atawneh and Aljehani, 2023) focused on different deep learning architectures (CNNs, LSTMs, RNNs, BERT) to identify the emails as legitimate or phishing by analyzing different features of emails. The proposed model using BERT and LSTM, reaching an impressive accuracy of 99.61% on publicly available 4 different datasets, outperforming other methods. Nowadays, other variants of the BERT model such as DistilBERT, RoBERTa, and TinyBER (Jiao et al., 2019) are performing well in the phishing and spam detection research sector (Jamal et al., 2023; Lee et al., 2020; Songailaitė et al., 2023). The authors (Jamal et al., 2023) presented an Improved Phishing and Spam Detection Model (IPSDM) which consists of two BERT variants such as the fine-tuned DistilBERT and RoBERTA model and showed good accuracy in this research scope. Lee et al. Lee et al. (2020) proposed Context-Aware Tiny Bert (CatBERT) which exhibits faster processing and robustness to adversarial attacks than other transformer models, especially those that involve intentional keyword substitutions with synonyms or typos. The detection rate of this proposed model was 87% which was far better than other compared models. Furthermore, researchers also used fine-tuned BERT-based models for phishing URL detection (Wang et al., 2023; Maneriker et al., 2021).
Overall, this literature survey shows the existing works although showing good performance by using different datasets and models but most of the research didn’t work model’s interpretability or explainability. “How does a model work?” or “Which things or materials contribute to detecting phishing?” those types of questions can arise in their working procedure. In this work, to answer those types of questions and solve the existing drawbacks in the literature, we have worked with explainable AI using not only one technique but also used two different techniques with different methodologies. In the literature, although some researchers did full experiments for phishing email detection but they utilized spam datasets as phishing datasets. In cybersecurity, we know that two different types of emails exist: spam and phishing, and not all spam emails are considered to be phishing. For this reason, we only focused on the phishing email dataset.
3 Methodology
In this paper, we have tried to explore the Transformer-based self-attention mechanism BERT model with the explainable AI. Initially, we collected a dataset and prepared it with data cleaning and balancing techniques. After preparing the data we have focused on optimizing and fine-tuning the model. For this phishing email detection purpose, we have selected the DistilBERT model and pre-trained with the prepared data. The DistilBERT model demonstrates good performance in both balanced and imbalanced scenarios of datasets when identifying phishing emails from texts. At the end of the experiment, we employed two XAI techniques Explanations LIME and Transformer Interpret to explore our fine-tuned model and described how it is working and performing with good accuracy in this text classification. The overall methodology of the proposed method is depicted in Fig. 1.
Figure 1: The proposed overall methodology.
3.1 Data Collection and Processing
For this experiment, a publicly available dataset has been collected from kaggle. The dataset describes different kinds of email text bodies that can be used to identify phishing emails using in-depth text analysis. This dataset consists of “Safe Email” and “Phishing Email”. Fig. 2 shows the data overview. The dataset has 11,322 Safe Emails and 7,328 Phishing Emails which shows that it is an imbalanced dataset. Along with this, this dataset showed that null values exist in rows. As null values can’t be processed that’s why those rows are removed. By using the resampling technique minor class “Phishing Email” is oversampled. This resampling technique mainly generates copies of other instances and adds those copies to the dataset. It is essential to address uneven data in order to develop equitable and precise predictive models. Balancing data ensures algorithms work well across all categories by reducing biases toward the majority class. This prevents skewed outputs and makes it possible to create a more accurate and equitable model for underrepresented classes. After oversampling, the dataset has 11,322 safe emails and 11,322 phishing emails in equal proportion. This data balancing also helps to reduce the risk of oversampling. Fig. 3 illustrates the imbalance and balanced data before and after sampling.
Figure 2: The sample data overview
(a) Imbalanced data
(b) Balanced data
Figure 3: Imbalanced and balanced data.
3.2 Dataset splitting
For this experiment, we have split the dataset into 70-30% where 70% data is training data and 30% is testing data. In 70% training set total of 15,850 data exist and the testing set consists of 6,794 data. Initially, the training set was utilized for the train the model, and the testing test was used to find out the effectiveness of the model on unseen data identification and prediction is phishing or not.
3.3 Model Selection
In this experiment, we have employed the transformer-based BERT variant architecture DistilBERT which is a lightweight, fast, and economical model. This is developed for language processing tasks and also compresses the original BERT model and makes it faster. Knowledge distillation was utilized in the pre-training phase of this pre-trained model, which allowed it to be effectively reduced in size by 40% while maintaining 97% of its language understanding skills and delivering a 60% speed increase (Sanh et al., 2019). This small variant has twelve attention heads, 768 hidden layers, and six transformer layers (Adoma et al., 2020). Whereas DistilBERT contains about 66 million parameters, the original BERT model has 110M (million) parameters. By lowering the number of parameters, DistilBERT maintains excellent performance while achieving a smaller size and faster inference speed.
3.3.1 Model Optimization
To improve its handling of terms that are not in its lexicon and to capture meaningful sub-word representations, BERT uses a Word-Piece tokenizer, which splits words into smaller sub-words. The preprocessed phishing dataset is tokenized using the tokenizer of the DistilBERT transformer. In this tokenized portion, the maximum sequence length of input data has been used 512. Other crucial components of tokenization are padding and truncation. The major purpose of these procedures is to ensure that each tokenized input sequence has a constant length. This is important because it creates consistent embeddings and efficient processing in batches.
Pre-trained weights acquired throughout the pre-training phase are used to initialize the DistilBERT models. Selecting an optimal batch size is paramount in deep learning, as it significantly influences both training and testing accuracies, along with overall runtime (Lin, 2022). The careful selection of batch size determines the efficiency of the neural network training process, striking a balance between computational resources and model performance. The batch sizes for training and testing data are optimized at 32 and 64, respectively. Each epoch’s training data is assigned at random to ensure the model sees a diversity of unseen data samples and prevents overfitting. This shuffling method increases the model’s ability to generalize.
During training, an optimizer is essential in adjusting the parameters of a neural network. The main goal of this is to minimize the error or loss function of the model for better performance. For this research work, AdamW (Adam Weight Decay) (Loshchilov and Hutter, 2017) optimizer has been used to get better performance and avoid overfitting. This optimizer is a modification version of Adam which is widely adopted for training deep neural networks (Zhuang et al., 2022). By separating weight decay from the learning rate, AdamW enhances its implementation. It improves regularization and training dynamics by separating weight decay from gradient updates, in contrast to Adam. Learning rate is a crucial hyperparameter governing step size during training for optimizing models. The convergence and overall performance of the system are dependent on various aspects such as the model architecture, optimization techniques, and task domain. For efficient fine-tuning and fruitful model training, this parameter must be adjusted. Selecting a high learning rate can introduce instability, resulting in poor performance on unseen data, while a low learning rate may extend the training process, requiring more epochs and increasing computational costs. In this research, a learning rate, 2e-5 is used in the training process for the transformer-based DistilBERT model.
The overall discussion about model optimization indicates the parameter settings for the introduced best model for this specific experiment. The hyperparameter settings for the suggested model are shown in Table 1. We worked with every possible value of the parameters by using an iterative process to find the ideal value. Following that, these improved hyperparameter settings are used to train the suggested model.
Table 1: Parameter Settings for Phishing Email Detection Task.
Parameter Name Worked With Optimal Value
Training Batch size 8, 16, 32 32
Testing Batch size 8, 16, 32, 64 64
Optimizer AdamW AdamW
Learning rate 1e-5, 2e-5, 3e-5 2e-5
Number of epochs 3, 4, 5, 6, 7 6
3.3.2 Fine Tuning
A pre-trained model is fine-tuned using task-specific data and backpropagation to alter parameters on certain tasks and datasets. The primary objective of this is to improve the model’s performance across various tasks and datasets while minimizing the loss function. In this experiment, DistilBERTForSequenceClassification has been utilized to adopt the pre-trained model. DataLoader is used to configure the train_loader and test_loader by setting prepared data, batch size, and shuffling command where train_loader is configured with shuffling and test_loader is configured without shuffling. After that, we fine-tuned this DistilBERT model by adjusting and adopting some necessary steps which are shown in Fig.  4.
Figure 4: The overall process of fine-tuning.
In this overall fine-tuning process, following the optimizer’s setup, the model is fine-tuned by feeding it the input parameters (input_ids, attention_mask, and label) during the forward pass so that predictions can be made. After the forward pass, gradients of the loss for the parameters (weights and biases) are calculated in the backward pass. These gradients show how changing the parameters would affect the loss. Lastly, an update is made to the model parameters. The optimizer is mainly responsible for changing the neural network’s parameters in accordance with the calculated gradients. It adjusts the weights and biases in an attempt to reduce the loss. The model gains the capacity to fine-tune its parameters throughout several epochs, which improves its predictive capability on the training set. The ultimate goal is to identify underlying patterns in the training set and use those patterns to generalize well to new data. Algorithm 1 shows the overall fine-tuning process in training.
Algorithm 1 Fine-Tuning the DistilBERT model
1:DistilBERT model, training data loader, optimizer, and number of epochs
2:Trained fine-tuned model
3:$\text{epochs}\Leftarrow n$
[…]
5:for epoch in range(epochsdo
6:     model.train()
7:     for batch in train_loader do
[…]
10:         $\text{labels}\Leftarrow\text{batch}[2]$
11:         $\text{gradients}\Leftarrow\text{0}$
[…]
13:         loss_calculation()
14:         backwardpass()
15:         update_parameters()
16:     end for
17:end for
3.3.3 Model Explainability
The ability of a model to explain its decisions is crucial since it provides insight into the process. Verifying the accuracy of judgments and reducing errors are made feasible by understanding how models work. Explainable AI has been a popular area with a growing group that has developed efficient techniques to understand deep neural network predictions among the more complex machine learning models (Holzinger et al., 2022). This is important because sophisticated models frequently act as opaque “black boxes" making it difficult for people to comprehend the logic behind particular decisions. In this research work, we have worked with model explanability to solve the black box problem. To understand the model prediction two explainability models LIME and Transformers Interpret have been utilized.
Generally, this experiment for phishing email detection is a text classification task that LimeTextExplainer used to explain the black box DistilBERT model. The purpose of the LIME method is to faithfully explain any classifier or regressor’s predictions (Ribeiro et al., 2016). For improved interpretability, LIME provides an understandable depiction of the decision-making process, providing light on the “why" behind the predictions or suggestions made by these models. In order to achieve model explainability, LIME first perturbs the email content by making modest adjustments, rearranging words, or adding new words. After that, the black-box transformer-based model is applied to the altered email content to generate predictions. From these impacted instances, LIME derives features. This local model should roughly represent the behaviour of the black-box model surrounding the chosen instance. Next, explanations for the black-box model’s prediction for the original data are produced using the local model. This explanation usually highlights the words or key elements most important during the model’s decision-making.
A Transformers Interpret tool was also utilized to explain the transformer-based model for this classification work. The class explainer process was carried out using this tool using a model and tokenizer. With the use of this class explanation, word attributions are computed. Positive attribution numbers signify a word’s positive contribution to the projected class, while negative numbers signify a word’s negative contribution. The terms that demonstrate a strong contribution to decision-making are also highlighted in this explanation. The overall process of model explainability is shown in Fig. 5. Algorithm 2 shows the explanation process of the model prediction.
Figure 5: The overall process of model explainability
Algorithm 2 Explainability Analysis
1:Trained model, tokenizer, and text
2:Explanation of the model prediction
3:class_names $\Leftarrow$ [X, Y, …N]
4:function Explain_Prediction(model, tokenizer, text)
5:     LIME_Explanation(model, text)
6:     Transformer_Interpret(model, tokenizer, text)
7:end function
8:function LIME_Explanation(model, text)
9:     explainer $\Leftarrow$ LimeTextExplainer(class_names $\Leftarrow class\_names$)
10:     explanation $\Leftarrow$ explainer.explain_instance(text, model, num_features $\Leftarrow n$, top_labels $\Leftarrow 1$)
11:     exp_class $\Leftarrow$ explanation.available_labels()
12:     exp $\Leftarrow$ explanation.as_list($\Leftarrow$ exp_class)
13:     explanation_visualize()
14:end function
15:function Transformer_Interpret(model, tokenizer, text)
16:     cls_explainer $\Leftarrow$ SequenceClassificationExplainer(model, tokenizer)
17:     word_attributions $\Leftarrow$ cls_explainer(text)
18:     cls_explainer_visualize()
19:end function
4 Results
The fine-tuned transformer-based model was trained and tested with both unbalanced and balanced datasets. To assess the performance of the model different measurements such as confusion matrix, precision, recall, F1-score, and overall accuracy have been calculated. Overall summarize the performance, a model needs to calculate the confusion matrix first. It is a kind of table that provides a detailed breakdown of the actual and predicted classification for specific task domain data. The confusion matrix of the DistilBERT model for both imbalanced and balanced datasets is shown in Table 2. True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN) are the main components of the confusion matrix and as well as are needed for the calculation of Precision, recall, and F1-score (Uddin et al., 2022). The precision, recall, and F1-score of the DistilBERT model for both imbalanced and balanced datasets are shown in Table 3 and Table 4.
Table 2: Confusion matrix of DistilBERT for both imbalanced and balanced dataset
Imbalanced Balanced
Class Name Safe Email Phishing Email Safe Email Phishing Email
Safe Email 3235 116 3281 98
Phishing Email 24 2216 5 3410
Table 3: The precision, recall, and f1-score for imbalanced dataset.
Class Precision Recall F1-score Support
Safe Email 0.99 0.97 0.98 3351
Phishing Email 0.95 0.99 0.97 2240
Table 4: The precision, recall, and f1-score for balanced dataset.
Class Precision Recall F1-score Support
Safe Email 1.00 0.97 0.98 3379
Phishing Email 0.97 1.00 0.99 3415
Precision: Precision which expresses the ratio of true positives to all predicted positives, indicates how accurate positive predictions are and emphasizes the model’s capacity to identify positive samples.
Precision $=\frac{TP}{TP+FP}$ (1)
Recall: It calculates the ratio of true positives to all predicted positives to assess a model’s accuracy in identifying positive samples. It denotes the sensitivity of the model, capturing the actual positive rate in the class observations.
Recall $=\frac{TP}{TP+FN}$ (2)
F1-score: A balanced statistical metric that assesses a model’s efficaciousness in predicting positive instances while accounting for imbalanced datasets is the F1-score, which is derived from the harmonic mean of accuracy and recall. It offers a thorough perspective that balances recall and precision, impacting the final score determined by the model’s projections of the positive rate and actual positive cases.
F1-score $=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (3)
The frequency of valid predictions in a model is quantified by accuracy, which also represents the percentage of correctly predicted samples out of all the samples the model generates.
$\text{Accuracy}=\frac{\text{TN}+\text{TP}}{\text{TN}+\text{TP}+\text{FN}+\text{FP}}$ (4)
The accuracy of a model can differ based on the various parameters, data balanced or not. Initially in this experiment, an imbalanced dataset was utilized to assess the model’s performance in both training and testing. This dataset is highly imbalanced based on the safe email and phishing email data where the number of data of safe email is quite higher than phishing email. To train the model with 70% imbalanced data we have used 6 epochs with a batch size of 32. After 6 epochs the model showed 98.90% training accuracy for this imbalanced dataset with 0.0201 training loss. On the other hand, we get a testing accuracy 97.50 with testing loss 0.0900 with batch size 64. Fig.  6 shows a graph of accuracy and loss with 6 epochs for an imbalanced dataset.
Figure 6: Training vs. Testing Accuracy and Loss on Imbalanced Data
Moreover, we also conducted this experiment with a resampled and balanced dataset. After preparing the balanced dataset again used 70% of the data for training purposes. Using this portion of data, we worked on model optimization and fine-tuning and trained the model. This time with balanced data the fine-tuned model showed 99.07% training accuracy with 0.0218 training loss. After that, we tested this trained model with 30% unseen balanced data and got 98.48% testing accuracy with 0.0447 testing loss. This performance showed that the performance of the model with a balanced dataset is quite better than an imbalanced dataset. Fig. 7 shows a graph of accuracy and loss with 6 epochs for a balanced dataset. The performance comparison of the model on both balanced and imbalanced dataset shows in Table 5.
Figure 7: Training vs. Testing Accuracy and Loss on Imbalanced Data.
Table 5: The Performance comparison of fine-tuned DistilBERT model on Imbalanced and Balanced dataset.
Dataset Mode
Training Loss
Training Accuracy
Testing Loss
Testing Accuracy
Imbalanced
0.0201
98.90
0.0900
97.50
Balanced
0.0218
99.07
0.0447
98.48
Finally, we have compared our Phishing Email Detection study with some previous studies in the same field in order to assess its quality. This comparison study is summarized in Table 6, which evaluates our work from several angles. These include the models used for machine and deep learning, the datasets used, and whether or not the study offers human interpretation and decision explainability for the model.
Table 6: Explainability of Different Studies on Phishing Email Detection
Reference
Datasets
Models
Explainability Analysis
2023(Jamal et al., 2023)
Email Spam Detection Dataset, PhishingEmailData
DistilBERT, RoBERTA
No
2021(Alhogail and Alsabih, 2021)
Fraud Dataset
GCN
No
2019(Fang et al., 2019)
IWSPA-AP
THEMIS
No
2023(Thapa et al., 2023)
IWSPA-AP, Nazario, Enron, CSIRO, Phisbowl
THEMIS, BERT
No
2023(Atawneh and Aljehani, 2023)
Phishing Email Dataset, Benign Email Dataset
BERT-LSTM
No
This Study Phishing Email Fine-tuned DistilBERT Yes
The comparison showed that our research work provides transparency in this phishing email detection task. However, some listed research work showed higher accuracy than our proposed model. This happened because they used different datasets in their experiment. In this study, we only focused on the phishing email dataset but other existing works were also done this work with the spam dataset. Furthermore, to the best of our knowledge, none of the papers listed in Table  6 offer a thorough human explanation of the classifier’s performance or clarify how the model functions in phishing email detection.
5 Explainability Analysis
LIME and Transformer Interpret both showed how the fine-tuned DistilBERT model arrives at its prediction on phishing email detection by approximating the model’s behaviour on every word of the email text. Lime and Transformer Interpret provided accurate results by answering the “Why” model predicting the selected text is a “Safe Email” or “Phishing Email”. Based on the explainability, those explainable techniques generate word attributions by providing positive and negative numbers on each word. Those numbers showed which word was more specifically reasonable for the predicted class. By using those numbers the explainable model visualizes the whole text by highlighting the words that scored positive and as well as negative on the prediction.
For this explainability analysis in LIME explainer, we initialized the number of features (words) 15 for the dimensionality reduction. If the number of words of a selected test is very large that time the computational time will increase as well as it increases the dimensionality which also can be a reason why the systems crashed. Limiting this number of features makes the explanation process computationally more feasible and provides a clearer and more concise picture of which words in the case of text are influential in the model’s prediction for a particular instance. Fig. 8 shows the model explainability using LIME and Table  7 and Table  8 show the positive and negative weights associated with words that help users understand phishing email detection.
(a) Explanation for the first text
(b) Explanation for the second text
Figure 8: LIME Explanation for two Phishing Email Texts
Table 7: The positive and negative weights using LIME for first text
Positive Numbers Negative Numbers
Word Value Word Value
‘low’ 1.760212017947662 ‘neonate’ -0.1267345497528774
‘price’ 1.5916451492409904 ‘3107’ -0.013956063314091143
‘software’ 1.2586059555064724
‘setupmefree’ 0.9966181271429526
‘com’ 0.3984979685547945
‘http’ 0.1447738056917774
‘a’ 0.06584106490791478
Table 8: The positive and negative weights using LIME for second text
Positive Numbers Negative Numbers
Word Value Word Value
‘megapixel’ 0.45124215111245936 ‘a’ -0.09469945665856638
‘get’ 0.4369880334209157
‘free’ 0.4206519368628723
‘digital’ 0.3721259726068145
‘camera’ 0.2388125335048898
‘canon’ 0.19498844421129996
‘6’ 0.04953986530836751
‘3’ 0.024635744735869915
Along with LIME, we also examined the model interpretability with similar text by using Transformer Interpret. During this interpretation, this Transformer Interpret uses the tokenizer to split the large words into sub-words with the ‘##’ prefix. This happens when an unusual or non-vocabulary word is encountered by the model, which breaks it down into smaller sub-words to enable efficient processing. Fig. 9 shows Transformer Interpretation for a phishing email and Table  9 and Table  10 show the word importance, with positive and negative numbers associated with each word for two phishing emails.
(a) Explanation for first text
(b) Explanation for second text
Figure 9: Transformer Interpret Explanation for two Phishing Email Text
Table 9: The positive and negative numbers using Transformer Interpret for first Phishing Email Text
Positive Numbers Negative Numbers
Word Value Word Value
‘low’ 0.7698059423920272 ‘.’ -0.09537831853568463
‘price’ 0.4465021887746993 ‘setup’ -0.037898186884516454
‘software’ 0.10363795047092349 ‘?’ -0.04837875144898304
‘http’ 0.05008396373925551 ‘a’ -0.212114259929307
‘:’ 0.06724251662885174 ’=’ -0.13427331689723981
‘/’ 0.018869178109329452 ‘310’ -0.06162690120821821
‘neon’ 0.07499325538696087 ‘##7’ -0.013360917838605864
‘##ate’ 0.09833698869154901 ‘.’ -0.1269640647507237
‘me’ 0.09196861346171359
‘free’ 0.16949987141093556
‘com’ 0.20340796716052334
‘/’ 0.006729002810807768
Table 10: The positive and negative numbers using Transformer Interpret for second Phishing Email Text
Positive Numbers Negative Numbers
Word Value Word Value
‘get’ 0.49733595483460524 ‘a’ -0.2688505206057748
‘free’ 0.4645088664478948 ‘6’ -0.09532782298942542
‘mega’ 0.18191195439136695 ‘.’ -0.08192496249808812
‘3’ 0.014327378740528552 ‘##pi’ -0.11609630028309895
‘##x’ 0.5855798135703232 ‘!’ -0.18838964048006548
‘digital’ 0.021290989158880363
‘canon’ 0.13188078942699152
‘camera’ 0.05155301993306188
‘>’ 0.03909299350598931
In Fig. 9 the highlighted text with green color shows the word importance to classify the analyzed text as Phishing. On the other hand, the red portions showed those words or sub-words that worked against the classified class. As we worked with a binary classification task here “True Label" = 1 means "Phishing Email" and when “True Label" = 0 means “Safe Email”.
According to the analysis, both explainable approaches explain how the suggested model predicts "Phishing Email" for each of the two tested input texts, but they differ in how they define explanation. LIME mostly uses small textual modifications to evaluate the impact of the model’s prediction on the entire text, while Transformer Interpret employs an attention mechanism to examine each segment of the input sequence. Furthermore, in the LIME and Transformer Interpret explanation processes, the resulting quantity of tokens, words, or sub-words differs for the two phishing email contents. The positive words are mainly responsible for the predicting class. More specifically, when those words are present in a text that time model predicts a class. A comprehensive difference between the two explainable AI techniques is shown in Fig. 10 and Fig. 11. The number of scoring percentages for those email texts is not the same for LIME and Transformer Interpret. This is only because the working procedures of those techniques are not the same.
(a) LIME Scoring
(b) Transformer Interpret Scoring
Figure 10: LIME and Transformer Interpret Comparison on first text
(a) LIME Scoring
(b) Transformer Interpret Scoring
Figure 11: LIME and Transformer Interpret Comparison on second text
6 Discussion
The results of both an imbalanced and balanced dataset depict an enhancement in the performance of the proposed transformer-based fine-tuned DistilBERT model. The fine-tuned version of the DistilBERT model showed 98.90% training and 97.50% testing accuracy on an imbalanced dataset whereas this model showed 99.07% training and 98.48% testing accuracy on the balanced dataset. Although these results show that the model’s performance on a balanced dataset is quite better than an imbalanced dataset, the difference is not so high, approximately 1%. This demonstrated that the fine-tuned version performs well in an imbalanced dataset. Additionally, the performance shows that the optimized model performed better than other previous studies.
Besides, this research work also answered the “How was the model performed?" and “Why was it performed?" using XAI, motivated from recently published book by Sarker (Sarker, 2024). Two well-known Explainable and interpretable techniques are utilized for the model understanding. It is more difficult to understand how DistilBERT makes its predictions because, as we know, it is a complex model. These AI techniques aided in gaining an understanding of the model’s decision-making process. Both LIME and Transformer Interpret highlighted the areas of the text that are more responsible for the classification of the phishing of safe email. This highlights areas worked with words or sub-words, making the decision more understandable for humans, and users can also comprehend it easily. Mainly, those techniques calculated each word or sub-words numeric value or weight where some words indicate positive value and some negative. Words with calculated higher positive values are more important for classifying the predicted class, while those with more negative values do not support the predicted class. This experimental analysis showed the transparency of the model in this phishing email classification task.
7 Conclusion
This study has investigated the detection of phishing emails by utilizing a Large Language Model. This also has shown how effective it is to use state-of-the-art natural language processing methods to improve the precision and effectiveness of phishing email identification. The overall experiment was done with both a balanced and imbalanced dataset. We have obtained promising results in differentiating between phishing and valid emails by utilizing DistilBERT’s contextual language understanding technique. Mainly, to achieve this good performance the model is optimized by changing some important parameters such as batch size, optimization function, learning rate, etc, and also fine-tuning has been done. This fine-tuned version of DistilBERT obtained 98.48% testing accuracy in this phishing email classification task. Furthermore, LIME and Transformer Interpret techniques have been employed to see the explanation of the model and how it makes decisions. The decision-making process is now more visible and understandable for users because of the usage of explainable AI, which promotes comprehension and trust. The overall model explainability describes the robustness of the proposed model.
The field of Large Language Models has experienced quick and dynamic development due to the substantial investment made by both businesses and consumers. New open-source models are being produced almost every day, demonstrating the ongoing development and innovation in this field. In the future, we plan to expand the scope of our research by experimenting with a wider variety of models and datasets. Furthermore, we intend to investigate the use of different variants of BERT-based models for the identification of different language phishing emails. Moreover, deep analysis of the model will be conducted to ensure its robustness by enhancing the explainable AI techniques.
[…]
Data Availability Statement
The dataset that is used in this study is available at Kaggle (Online Link):  https://www.kaggle.com/datasets/subhajournal/phishingemails.